# Covariances for Free: Exploiting Mean Distributions for Federated Learning with Pre-Trained Models

### v1

@@ 5 Experiments
Table 2: Evaluation of different training-free methods using 100 clients for four datasets and 9275 pre-defined clients on iNat-120K using 5 random seeds. We show the total communication cost (in MB) from all clients to server. We also show the FedCOF oracle in which full class covariances are shared from clients to server. The best results from each section are highlighted in bold.
@@
CIFAR100 FedNCM (Legate et al., 2023a) 41.5$\pm$0.1 5.9 55.6$\pm$0.1 14.8 55.2$\pm$0.1 8.9
Fed3R (Fanì et al., 2024) 56.9$\pm$0.1 110.2 62.7$\pm$0.1 670.1 73.9$\pm$0.1 244.8
FedCOF (Ours) 56.1$\pm$0.2 5.9 63.5$\pm$0.1 14.8 73.2$\pm$0.1 8.9
FedCOF Oracle (Full Covs) 56.4$\pm$0.1 3015.3 63.9$\pm$0.1 18823.5 73.8$\pm$0.1 6780.0
IN-R FedNCM (Legate et al., 2023a) 23.8$\pm$0.1 7.1 37.6$\pm$0.2 17.8 32.3$\pm$0.1 10.7
Fed3R (Fanì et al., 2024) 37.6$\pm$0.2 111.9 46.0$\pm$0.3 673.1 51.9$\pm$0.2 246.6
FedCOF (Ours) 37.8$\pm$0.4 7.1 47.4$\pm$0.1 17.8 51.8$\pm$0.3 10.7
FedCOF Oracle (Full Covs) 38.2$\pm$0.1 3645.7 48.0$\pm$0.3 22758.8 52.7$\pm$0.1 8197.4
CUB200 FedNCM (Legate et al., 2023a) 37.8$\pm$0.3 4.8 58.3$\pm$0.3 12.0 75.7$\pm$0.1 7.2
Fed3R (Fanì et al., 2024) 50.4$\pm$0.3 109.6 58.6$\pm$0.2 667.3 77.7$\pm$0.1 243.1
FedCOF (Ours) 53.7$\pm$0.3 4.8 62.5$\pm$0.4 12.0 79.4$\pm$0.2 7.2
FedCOF Oracle (Full Covs) 54.4$\pm$0.1 2472.1 63.1$\pm$0.5 15432.7 79.6$\pm$0.2 5558.6
Cars FedNCM (Legate et al., 2023a) 19.8$\pm$0.2 5.4 30.0$\pm$0.1 13.5 26.2$\pm$0.4 8.1
Fed3R (Fanì et al., 2024) 39.9$\pm$0.2 110.2 41.6$\pm$0.1 668.8 47.9$\pm$0.3 244.0
FedCOF (Ours) 44.0$\pm$0.3 5.4 47.3$\pm$0.5 13.5 52.5$\pm$0.3 8.1
FedCOF Oracle (Full Covs) 44.6$\pm$0.1 2767.3 47.2$\pm$0.3 17275.7 53.1$\pm$0.1 6222.5
iNat-120K FedNCM (Legate et al., 2023a) 21.2$\pm$0.1 111.8 36.0$\pm$0.1 279.5 53.9$\pm$0.1 167.7
Fed3R (Fanì et al., 2024) 32.1$\pm$0.1 9837.3 41.5$\pm$0.1 61064.1 62.5$\pm$0.1 22050.2
FedCOF (Ours) 32.5$\pm$0.1 111.8 44.1$\pm$0.1 279.5 63.1$\pm$0.1 167.7
FedCOF Oracle (Full Covs) 32.4$\pm$0.1 57k 43.6$\pm$0.1 358k 62.9$\pm$0.1 128k
Datasets. We evaluate the proposed method on multiple datasets namely CIFAR-100 (Krizhevsky, 2009), ImageNet-R (Hendrycks et al., 2021), CUB200 (Wah et al., 2011), Stanford Cars (Krause et al., 2013) and iNaturalist (Van Horn et al., 2018). CIFAR-100 has 100 classes provided in 50k training and 10k testing images. ImageNet-R (IN-R) is composed of 30k images covering 200 ImageNet classes. CUB200 has as well 200 classes of different bird species provided in 5994 training and 5794 testing images. Stanford Cars has 196 classes of cars with 8144 training images and 8041 test images. We distribute these datasets to 100 clients using a highly heterogeneous Dirichlet distribution $(\alpha=0.1)$ following standard practice (Hsu et al., 2019; Legate et al., 2023a). We also use the real-world non-iid FL benchmark of iNaturalist-Users-120K (Hsu et al., 2020) (iNat-120K) having 1203 classes across 9275 clients and 120k training images.
Figure 5: Performance vs. communication cost analysis using pre-trained SqueezeNet and ViT-B/16 on iNaturalist-Users-120K.
Implementation Details. We use three models: namely SqueezeNet (Iandola et al., 2016) following Legate et al. (2023a) and Nguyen et al. (2023), MobileNetV2 (Sandler et al., 2018) following Fanì et al. (2024); Hsu et al. (2020), and ViT-B/16 (Dosovitskiy et al., 2021). All models are pre-trained on ImageNet-1k (Deng et al., 2009). We use the FLSim library for our experiments and implement all methods in the same setting. We use $\gamma=1$ for all experiments with SqueezeNet and ViT-B/16, and $\gamma=0.1$ for all experiments with MobileNetV2 due to very high dimensionality $d$ of the feature space. Following Fanì et al. (2024), we use $\lambda=0.01$ for both Fed3R and FedCOF. We also include the oracle setting of FedCOF in which the full class covariances are shared from clients to server and aggregated similar to Luo et al. (2021). We discuss this oracle further in Appendix G. We also evaluate how FedCOF impact further fine-tuning (FT) or linear probing (LP) of the models. For all FT and LP experiments, we train 100 client models for 200 rounds, 1 local epochs per round, and set the client participation to 30%. We provide more details in Appendix F.
5.1 Evaluation for different training-free methods
@@
We compare the performance of existing training-free methods and the proposed method in Table 2 using pre-trained Squeezenet, Mobilenetv2 and ViT-B/16 models. We observe that Fed3R (Fanì et al., 2024) using second-order statistics outperforms FedNCM (Legate et al., 2023a) by significant margins ranging from 0.3% to 21% across all datasets. However, Fed3R requires a higher communication cost compared to FedNCM. In real-world iNat-120K benchmark, Fed3R needs $61k$ MB compared to $280$ MB for FedNCM (see Fig. 5), which is 218 times higher. FedCOF performs better than Fed3R in most settings despite having the same communication cost as FedNCM. FedCOF achieves very similar performance as the oracle setting using aggregated class covariances requiring very high communication, which validates the effectiveness of the proposed method. Note that in Table 2 the number of means used to estimate each class covariance is less than the total number of clients, e.g. approximately 11.8 means per class are used for estimation in CUB200. This is due to the fact that not all classes are present in all clients.
@@
FedCOF maintains similar accuracy with Fed3R on CIFAR100 and ImageNet-R, with an improvement of about 1% when using MobileNetv2. FedCOF performs significantly better than Fed3R on CUB200 and Cars. On CUB200, FedCOF outperforms Fed3R by 3.3%, 3.9% and 2.2% using SqueezeNet, MobileNetv2 and ViT-B/16 respectively. FedCOF improves over Fed3R in the range of 4.1% to 5.7% on Cars. On iNat-120K, FedCOF improves over Fed3R by 0.4%, 2.6% and 0.6% using different models, and even performs marginally better than the oracle setting which can be attributed to less noisy covariance estimation from large number of client means. When comparing FedCOF with FedNCM – both with equal communication costs and same strategy in clients – one can observe that the usage of second order statistics derived only from the class means of clients leads to large performance gains, e.g. 24% using SqueezeNet and 26% using ViT-B/16 on Cars, about 10% using all architectures on large-scale iNat-120K.
Figure 7: Analysis of performance when initialized with different methods and then fine-tuned with FedAdam (Reddi et al., 2020) using a pre-trained SqueezeNet model. Here, FedAdam (in blue) uses random classifier initilization and pre-trained backbone.
Figure 9: Linear probing with FedAvg (McMahan et al., 2017) using pre-trained ViT-B/16.
5.2 Analysis of Fine-tuning and Linear Probing
@@
While we achieve very high accuracy without any training with FedCOF, we show in Fig. 7 that further fine-tuning the model with FL optimization methods achieves better and faster convergence compared to federated optimization from scratch. We use adaptive optimizer, FedAdam (Reddi et al., 2020) for all FT experiments since it performs better than other optimizers as discussed in (Nguyen et al., 2023). We show in Fig. 7, how the performance of FedNCM, Fed3R, and the proposed FedCOF evolves over multiple rounds. These training-free methods end after all clients appear at least once to share their local statistics to server. We observe that fine-tuning after performing classifier initialization with FedCOF starts with a higher accuracy and converges faster and better compared to FedNCM. While FedCOF and Fed3R initialization converges similarly on ImageNet-R, FedCOF+FedAdam achieves a better accuracy than Fed3R+FedAdam on CUB200 and Cars. All the training-free classifier initialization approaches outperform FedAdam with a random classifier initialization and pre-trained backbone.
@@
Following (Legate et al., 2023a; Nguyen et al., 2023), we also perform federated linear probing of the models using FedAvg (McMahan et al., 2017) after classifier initialization with different training-free methods. Linear probing requires much less computation compared to fine-tuning the entire network and were found to be effective with pre-trained models. We observe that linear probing after FedCOF improves significantly compared to FedNCM and Fed3R using ViT-B/16 Fig. 9 on Cars. We provide more experiments in Appendix H. The improved performance with FedCOF initialization for finetuning and probing validates the effectiveness of the proposed method.
5.3 Ablation Studies
Figure 11: Ablation experiments on CIFAR-100: (a) Shows that communicating multiple class means per client improves FedCOF performance with fewer clients. (b) Shows how the performance of the proposed method changes with the number of clients and varying data heterogeneity. (c) Performance comparison of FedCOF with full, diagonal, and spherical covariance matrix communication.
Communicating multiple class means per client. We analyze the FL settings with fewer clients ranging from 10 to 50 clients in Fig. 11(a) and show that sharing more than 1 class means from each client by subsampling data and computing multiple means in each client improves the performance. Using only 10 clients, sharing 2 class means per client improves the accuracy by 2.6%.
@@ 5.3 Ablation Studies
Impact of the number of clients and data heterogeneity. We analyze the impact of the number of clients and data heterogeneity on the performance of FedCOF using MobileNetv2 on CIFAR-100 as shown in Fig. 11(b). We observe that the performance of FedCOF improves with increasing number of clients and decreasing heterogeneity. This is due to the fact that more clients provides more class means and more uniform data distribution gives better representative local means. While more number of clients are favourable for FedCOF, it still performs well and outperforms FedNCM significantly in the setting with only 10 clients and high data heterogeneity.
Communicating diagonal or spherical covariances. While communicating diagonal or spherical covariances (mean of the diagonal covariance) from clients to server and then estimating the global class covariance from them can significantly reduce the communication cost, such estimates of global class covariance is poor compared to FedCOF. We show in Fig. 11(c) that FedCOF outperforms these covariance sharing baselines when communicating spherical or diagonal covariances.
@@ 6 Conclusion
In this work we proposed FedCOF, a novel training-free approach for federated learning with pre-trained models. By leveraging the statistical properties of distribution client class sample means, we show that second-order statistics can be estimated using only class means from clients, thus reducing communication costs. We derive a provably unbiased estimator of population class covariances, enabling accurate estimation of a global covariance matrix. After applying shrinkage to the estimated class covariances and removing between-class scatter matrices, we show that the server can effectively use this global covariance to initialize the global classifier. Our experimental results show that FedCOF outperforms FedNCM (Legate et al., 2023a) by significant margins while maintaining the same communication costs. Additionally, FedCOF delivers competitive or even superior results to Fed3R (Fanì et al., 2024) across various model architectures and benchmarks while substantially reducing communication costs. Moreover, we empirically demonstrate that our approach can serve as a more effective starting point for improving the convergence of standard federated fine-tuning and linear probing methods.
Limitations. The quality of our estimator depends on the number of clients, as shown in Fig. 11(a) where using multiple class means per client helps with fewer client settings. Another limitation is the assumption that samples of the same class are iid across clients, which is however an assumption underlying most of federated learning. Finally, our method assumes the existence of a pre-trained network. If the domain shift with the client data is sufficiently large, this is expected to impact the performance.
@@ Appendix A Scope and Summary of Notation
  • $N$: total number of samples.
@@
Appendix B Proof of Proposition 1
Proposition 1.
@@ Appendix B Proof of Proposition 1
is distributed with mean $\mathbb{E}[\overline{F}_{k,c}]=\mu_{c}$ and covariance $\mathrm{Var}[\overline{F}_{k,c}]=\frac{\Sigma_{c}}{n_{k,c}}$.
@@
where in the last equality we used Eq. 13. Thus the expectation of the sample mean is $\mu$, which completes the first part of the proof.
@@
By the independence assumption of $\{F^{j}_{k}\}_{j=1}^{n_{k}}$, the cross terms $\mathrm{Cov}[F_{k}^{i},F_{k}^{j}]=0$ for $i\neq j$. Applying Equation 13, we have:

### v3

@@ 5 Experiments
Datasets. We evaluate FedCOF on multiple datasets namely CIFAR-100 (krizhevsky2009learning, ), ImageNet-R (hendrycks2021many, ) (IN-R), CUB200 (wah2011caltech, ), Stanford Cars (krause20133d, ) and iNaturalist (van2018inaturalist, ). We distribute the first 4 datasets to 100 clients using a highly heterogeneous Dirichlet distribution $(\alpha=0.1)$ following standard practice (hsu2019measuring, ; legate2023guiding, ). We also use real-world non-iid FL benchmark of iNaturalist-Users-120K (hsu2020federated, ) (iNat-120K) having 1203 classes across 9275 clients. We discuss the dataset details in Appendix˜J.
Implementation details. We use three models: namely SqueezeNet (iandola2016squeezenet, ) following legate2023guiding and nguyen2022begin , MobileNetV2 (sandler2018mobilenetv2, ) following fani2024accelerating ; hsu2020federated , and ViT-B/16 (dosovitskiy2021an, ). All models are pre-trained on ImageNet-1k (deng2009imagenet, ). We use the FLSim library. We use $\gamma=1$ for all experiments with SqueezeNet and ViT-B/16, and $\gamma=0.1$ for all experiments with MobileNetV2 due to very high dimensionality $d$ of the feature space. We compare to FedCOF Oracle in which real class covariances are shared from clients and aggregated in server instead of using our estimated covariances (see Appendix˜J). For all experiments, we set the client participation in each round to 30%, and we show the training-free methods in multiple rounds in Figures˜3 and 4. We provide more implementation details in Appendix˜J. We discuss computation of communication costs for all methods in Appendix˜H.
Table 3: Evaluation of different training-free methods using 100 clients for first four datasets and 9275 pre-defined clients on iNat-120K across 5 random seeds. We show the total communication cost (in MB) from all clients to server. We also show the FedCOF oracle in which full class covariances are shared from clients to server. Best results from each section in bold.
@@
CIFAR-100
[…]
IN-R
[…]
CUB200
[…]
Cars
[…]
iNat-120K
[…]
@@
Evaluation for different training-free methods. We compare the performance of existing training-free methods and the proposed method in Table˜3 using pre-trained Squeezenet, Mobilenetv2 and ViT-B/16 models. We observe that Fed3R (fani2024accelerating, ) using second-order statistics outperforms FedNCM (legate2023guiding, ) significantly ranging from 0.3% to 21% across all datasets. However, Fed3R requires a higher communication cost compared to FedNCM. In real-world iNat-120K benchmark, Fed3R needs $61k$ MB compared to $280$ MB for FedNCM (see Figure˜1), which is 218 times higher. FedCOF performs better than Fed3R in most settings despite having the same communication cost as FedNCM. FedCOF achieves similar performance as the oracle setting using aggregated class covariances requiring very high communication, which validates the effectiveness of the proposed covariance estimator.
@@
FedCOF maintains similar accuracy with Fed3R on CIFAR-100 and ImageNet-R, with an improvement of about 1% when using MobileNetv2. FedCOF outperforms Fed3R on CUB200 and Cars. On CUB200, FedCOF outperforms Fed3R by 3.3%, 3.9% and 2.2% using SqueezeNet, MobileNetv2 and ViT-B/16 respectively. FedCOF improves over Fed3R in the range of 4.1% to 5.7% on Cars. On iNat-120K, FedCOF improves over Fed3R by 0.4%, 2.6% and 0.6% using different models. When comparing FedCOF with FedNCM – both with equal communication costs and same strategy in clients – one can observe that the usage of second order statistics derived only from the class means of clients leads to large performance gains, e.g. 24% using SqueezeNet and 26% using ViT-B/16 on Cars, above 8% using all architectures on large-scale iNat-120K. We also adapt CCVR (luo2021no, ) for classifier initialization in Appendix˜K and show that FedCOF outperforms CCVR.
Comparison with communication-efficient methods. We consider the recent Probabilistic Federated Prompt-Tuning (PFPT) approach that aggregates trainable prompt parameters from clients and tunes them in a federated manner while keeping the backbone fixed (weng2024probabilistic, ). We compare PFPT, FedAvg-PT, and FedProx-PT (prompt-tuning variants of FedAvg and FedProx) with the proposed FedCOF. Similar to PFPT (weng2024probabilistic, ), we use a pre-trained ViT-B/32, assign class samples to clients using a Dirichlet distribution ($\alpha=0.1$), and use 3 random seeds. We use the same training hyperparameters as PFPT. We show in Table˜4 that FedCOF is much more communication efficient compared to PFPT and achieves higher performance without any training. We also observe that prompt-tuning methods perform poorly on fine-grained datasets. We discuss more on this in Appendix˜H.
Table 4: Comparison of FedCOF with federated prompt-tuning methods using ViT-B/32.
CIFAR-100 IN-R CUB200 Cars
Method Acc ($\uparrow$) Comm. ($\downarrow$) Acc ($\uparrow$) Comm. ($\downarrow$) Acc ($\uparrow$) Comm. ($\downarrow$) Acc ($\uparrow$) Comm. ($\downarrow$)
[…]
Comparison with full fine-tuning methods. We compare training-free methods with FL baselines like FedAvg and FedAdam with randomly initialized classifier and pre-trained backbone in Table˜5. We use adaptive optimizer, FedAdam (reddi2020adaptive, ) since it performs better than most other optimizers as shown in nguyen2022begin . Without any training, FedCOF outperforms FedAvg in all settings and FedAdam by 7.3% on CUB200 and 2.2% on Cars, and achieves competitive performance in ImageNet-R. We show in Figure˜3 how FedCOF starts from a very high accuracy compared to FedAdam and further improves on fine-tuning. We provide more experiments with pre-trained ResNet-18 in Appendix˜K.
Refer to captionRefer to captionRefer to caption
Figure 3: Performance comparison when initializing with different methods and fine-tuning with FedAdam (reddi2020adaptive, ) and FedAvg (mcmahan2017communication, ). We also compare with FedAdam and FedAvg using a pre-trained backbone and random classifier initialization. The training-free initialization stages are shown as dotted lines and stars represents the start of fine-tuning stages. Accuracies are averaged over 3 seeds.
Table 5: Comparison with training-based FL baselines (FedAvg and FedAdam) using pre-trained SqueezeNet. For training-based methods, we consider 100 rounds of training for fair comparison and report accuracy of 3 random seeds.
Method Training ImageNet-R CUB200 Cars
[…]
@@
Analysis of fine-tuning and linear probing. While FedCOF achieves high accuracy without training, we show in Figure˜3 that further fine-tuning the model with FL achieves better and faster convergence compared to federated optimization from scratch. We show the performance of fine-tuning after training-free classifier initialization in Figure˜3. These training-free methods end after all clients appear at least once to share their local statistics. We fine-tune the models after FedCOF and Fed3R for 100 rounds since they achieve fast convergence, while we train for 200 rounds for FedAdam, FedAvg, and fine-tuning after FedNCM which takes longer to converge. Fine-tuning after FedCOF starts at a higher accuracy and converges faster compared to FedNCM. Although FedCOF and Fed3R converge similarly on ImageNet-R, FedCOF+FedAdam achieves better accuracy than Fed3R+FedAdam on CUB200 and Cars. We see in Table˜5 that all training-free approaches followed by fine-tuning outperform FedAdam and FedAvg with random classifier initialization.
@@
Following legate2023guiding and nguyen2022begin , we perform federated linear probing (LP) of the models using FedAvg after classifier initialization with training-free methods. In FedAvg-LP, we perform FedAvg and learn only the classifier weights of all client models. Linear probing requires much less computation compared to fine-tuning the entire model and were found to be effective with pre-trained models. We observe in Figure˜4 that linear probing after FedCOF improves significantly compared to FedNCM and Fed3R using ViT-B/16 on Cars and SqueezeNet on iNat-120K. On the real-world dataset iNat-120K, FedAvg-LP with random classifier initialization achieves 27.3% after 5000 rounds while FedCOF+FedAvg-Lp achieves 34% in less than 1000 rounds. We plot accuracy versus communication in Figure˜4 (center) to demonstrate the advantage of FedCOF over other methods. We discuss more in Appendix˜K.
@@
Impact of number of clients and data heterogeneity. We analyze in Figure˜5 (left), the performance of FedCOF with varying number of clients and data heterogeneity. We observe that the performance of FedCOF improves with increasing number of clients and decreasing heterogeneity. This is due to the fact that more clients provides more class means and more uniform data distribution gives better representative local means. While more clients are favourable for FedCOF, it still performs well and outperforms FedNCM significantly in the setting with 10 clients and high data heterogeneity.
Multiple class means per client. We analyze FL settings with fewer clients ranging from 10 to 50 in Figure˜5 (center) and show that sharing multiple class means from each client improves the accuracy. Using only 10 clients, sharing 2 class means per client improves the accuracy by 2.6%.
Impact of shrinkage. We show in Figure˜5 (right) that using shrinkage improves the covariance estimates thereby improving the accuracy. We observe that shrinkage consistently improves performance, especially when the number of sampled means is small. When the number of means is low relative to the feature dimension ($d=512$ in SqueezeNet), the covariance estimate becomes ill-conditioned, and shrinkage stabilizes it. However, as more class means are sampled per client or the number of clients increases, the benefit of shrinkage diminishes, since the total number of means approaches the feature dimension, making the estimate more stable. We present more ablation studies in Appendix˜L.
Refer to captionRefer to captionRefer to caption
Figure 4: Performance of different classifier initialization methods when linear-probing with FedAvg (mcmahan2017communication, ). FedAvg-LP (in blue) uses random classifier initialization and a pre-trained backbone. The training-free initialization stage is shown in dotted lines, stars represents the start of linear probing.
Refer to captionRefer to captionRefer to caption
Figure 5: Ablation experiments: (left) Change in performance with varying number of clients and data heterogeneity. (center) Sharing multiple class means per client improves FedCOF performance. (right) Impact of shrinkage with varying number of clients and sampled means per client.
@@ 6 Conclusion
In this work we proposed FedCOF, a novel training-free approach for federated learning with pre-trained models. By leveraging the statistical properties of client class sample means, we showed that second-order statistics can be estimated using only class means from clients, thus reducing communication costs. We derived a provably unbiased estimator of population class covariances, enabling accurate estimation of a global covariance matrix. By applying shrinkage to the estimated class covariances and removing between-class scatter matrices, the server can effectively use this global covariance to initialize a global classifier. Our experiments demonstrated that FedCOF outperforms FedNCM (legate2023guiding, ) by significant margins while maintaining the same communication cost. Additionally, FedCOF delivers competitive or even superior results to Fed3R (fani2024accelerating, ) across model architectures and benchmarks while substantially reducing communication costs. Finally, we showed that FedCOF outperforms federated prompt-tuning methods and serves as a more effective starting point for improving the convergence of federated fine-tuning and linear probing methods.
Limitations. The quality of our estimator depends on the number of clients, as shown in Figure˜5 where using multiple class means per client helps with fewer clients. Another limitation is the assumption that samples of the same class are iid across clients, which is, however, an assumption underlying most of federated learning. We discuss the bias in our estimator in non-iid settings in Appendix˜G.
Acknowledgements.
We acknowledge project PID2022-143257NB-I00, financed by MCIN/AEI/10.13039/501100011033 and ERDF/EU and FSE+, funding by the European Union ELLIOT project, and the Generalitat de Catalunya CERCA Program. Bartłomiej Twardowski acknowledges the grant RYC2021-032765-I and National Centre of Science (NCN, Poland) Grant No. 2023/51/D/ST6/02846. Kai Wang acknowledges the funding from Guangdong and Hong Kong Universities 1+1+1 Joint Research Collaboration Scheme and the start-up grant B01040000108 from CityU-DG. Andrew Bagdanov acknowledges financial support from the AVIOS project, financed by the Italian Ministry of Enterprises and Made in Italy (MiSE/MIMIT). We acknowledge all reviewers, especially anonymous reviewer (Md8N) whose constructive feedback restored our belief in the review system, which had been severely shaken in earlier review rounds.
Appendix
@@ Appendix A Scope and Summary of Notation
  • NN: total number of samples.
@@
Appendix B Expectation and Variance of the Sample Mean
@@
is distributed with mean and covariance given by:
$\mathbb{E}[\overline{F}_{k,c}]=\mu_{c}\qquad\mathrm{Var}[\overline{F}_{k,c}]=\frac{\Sigma_{c}}{n_{k,c}}$ (10)
@@
where in the last equality we used Equation˜11. Thus the expectation of the sample mean is $\mu$, which completes the first part of the proof.
@@
By the independence assumption of $\{F^{j}_{k}\}_{j=1}^{n_{k}}$, the cross terms $\mathrm{Cov}[F_{k}^{i},F_{k}^{j}]=0$ for $i\neq j$. Applying Equation˜11, we have: